Explain the equilibrating process in the simple Keynesian model.
Answer #1

Answer

In simple kenesian model the equilibrium level of employment and output is determined at a point where aggregate demand function intersects aggregate supply function.

Aggregate demand function AD= C+I

C= consumption

I= investment

Aggregate supply function AS= C+S

C= consumption

S= saving

Equilibrium condition is AD= AS

Ie. C+I=C+S

C is common in both sides of the equation

Therefore,

I=S

point. AS = CAS - Egmile Equilibrium (t I AD= Aggregate demand/AS National Income.